# Benchmarking Universal Machine Learning Interatomic Potentials for Supported Nanoparticles: Decoupling Energy Accuracy from Structural Exploration

### v1

@@ 2 Results
Finite-Temperature Dynamics. The dynamic behavior of supported nanoparticles is important for understanding the catalytic activity and stability. Several MD-based tasks are designed to examine uMLIPs. We first tested the mean-squared displacements of Cu atoms (MSDCu, defined in Equation 4) of Cu13 on three Al2O3 surfaces with short-time MD simulations, where ab initio MD (AIMD) trajectories are available for reference. The DP-UniAlCu, MACE-UniAlCu, MACE-OMAT, MatterSim-v1.0.0-1M, MatterSim-v1.0.0-5M, and DPA3-OMAT are tested. For each MLIP, 20 MD simulations starting from the same global minimum configuration but initialized from different velocities by Maxwell-Boltzmann distribution at 800 K are performed for 20 ps with a timestep of 2 fs. For AIMD, only one trajectory is performed with the same settings. The window-averaged MSDCu up to 5 ps are shown in Figure 4a-c. The vertical bars are the standard deviations of 20 MD simulations by MLIPs. The model accuracy can be examined by checking whether the AIMD results fall within the standard deviations of MLIPs. For $\gamma$-Al2O3(100), all models except for MACE-OMAT and DPA3-OMAT give MSDCu results that are close to the AIMD results, which exhibit smaller MSDCu values. For $\gamma$-Al2O3(110), all models give consistent results with AIMD. For $\alpha$-Al2O3(0001), all models except for two MatterSim models produce MSDCu results closer to AIMD, which show larger MSDCu values. The radial distribution functions (RDFs) of Cu-Al and Cu-O pairs are shown in Figure 4d-f and Figure 4g-i, respectively, which are averaged over snapshots every 0.1 ps from 0.2 ps to 20 ps. The RDFs of Cu-Al and Cu-O by DP-UniAlCu, MACE-UniAlCu and MACE-OMAT are in good agreement with the AIMD results, while other models give larger deviations. The RDF of Cu-Cu can be well reproduced by all models (Figure S7). Since the 20-ps-long MD simulations are too short to observe any diffusion of the entire nanoparticle, we performed 1-ns-long MD simulations for Cu13 on three $\alpha$-Al2O3(0001) surfaces with DP-UniAlCu, MACE-UniAlCu, MACE-OMAT, MatterSim-v1.0.0-1M, and DPA3-OMAT to investigate the long-time dynamic behavior. Five independent MD simulations are performed for each MLIP starting from the same global minimum configuration but initialized from different velocities by Maxwell-Boltzmann distribution at 800 K. The results of MSDCu up to 100 ps for Cu13 on three Al2O3 surfaces are shown in Figure 5a-c. The vertical bars are the standard deviations of five MD simulations by MLIPs. Since ns-long MD simulations are not readily accessible to AIMD, the benchmark here only served as a cross-validation across different MLIPs by comparing results against our DP-UniAlCu model. Overall, all MLIPs predict qualitatively consistent mobility trend of Cu atoms on three Al2O3 surfaces, where the MSDCu on $\gamma$-Al2O3(100) and $\alpha$-Al2O3(0001) are comparable and larger than $\gamma$-Al2O3(110). MatterSim-v1.0.0-1M gives slightly larger MSDCu for Cu13 on $\gamma$-Al2O3(100) and $\gamma$-Al2O3(110), and much larger MSDCu for $\alpha$-Al2O3(0001) compared to DP-UniAlCu. MACE-OMAT gives almost identical MSDCu for Cu13 on $\gamma$-Al2O3(110) as DP-UniAlCu while more deviations between them are observed for Cu13 on $\gamma$-Al2O3(100) and $\alpha$-Al2O3(0001), where MACE-OMAT gives smaller MSDCu values. By using the MACE-UniAlCu model, which is trained on the same dataset as DP-UniAlCu, the MSDCu results for $\gamma$-Al2O3(100) and $\alpha$-Al2O2(0001) become closer to those by DP-UniAlCu, however, a smaller MSDCu is observed for $\gamma$-Al2O3(110). DPA3-OMAT exhibits a consistent mobility trend as DP-UniAlCu but gives smaller MSDCu values for all three Al2O3 surfaces. To perform a more quantitative comparison, more simulations with a longer time are necessary, which can be challenging for uMLIPs. The computation performance is shown in Figure 5d, which is taken from the 100-ps simulation of an FCC Cu bulk with 2048 atoms. All models are performed on a single NVIDIA 80G A100 GPU with LAMMPSThompson2022ComputPhysCommun except for MatterSim-v1.0.0-1M with ASELarsen2017JPhysCondensMatter. Our DP-UniAlCu model is about two orders faster than other models, achieving the best balance between accuracy and efficiency here. The reduced efficiency of the uMLIPs arises from their large parameter counts, the overhead associated with message passing, and the substantial cost of tensor products when equivariant features are employed.Batatia2025NatMachIntell Though quantitative results for large size and time scales may not be obtained directly by uMLIPs, they are useful models to sample configuration spaces of interest, which can be distilled by cheaper models realizing long-time simulations.
3 Discussion and Conclusion
@@
In this work, we tested several uMLIPs on supported Cu nanoparticles on Al2O3 surfaces and found that they can be very useful in simulating supported nanoparticles even without any fine-tuning, including the global optimization of low-energy configurations and the MD at finite temperature.
@@
The uMLIPs examined in this work can produce reasonable results on the binding energies of small supported nanoparticles (Cu1-21) reported in our previous study,Xu2025ACSNano where MACE-OMAT shows an accuracy comparable to our domain-specific DP-UniAlCu model. In the global optimization of larger supported nanoparticles (Cu27, Cu38, Cu47, and Cu55), our DP-UniAlCu model behaves better than MACE-OMAT and Mattersim-v1.0.0-1M in finding low-energy minimum configurations. Interestingly, Mattersim-v1.0.0-1M, which shows larger deviations in binding energies, can find even more stable configurations than the other two models in some systems, indicating its exploration ability in the configuration space is not solely characterized by the error metrics on a preset dataset. Alongside the exploration ability of MatterSim-v1.0.0-1M, the non-uniform accuracy across different systems is observed, for instance, the binding energy accuracy in one nanoparticle size is better than another size on the same Al2O3 surface. This non-uniform accuracy may be due to the imbalance in the number of configurations from different systems, where the initial dataset is made up of configurations from one specific system and the configurations of other systems are accumulated in active learning. In terms of global optimization, since the system size of interest is usually accessible to DFT, the MLIP-generated configurations can still be verified by DFT to find the global minimum configurations. The model’s stability in generating diverse configurations in a large region of the configuration space is more important than the accuracy of the energy, as it provides more candidates for DFT verification. A model with better accuracy can improve the performance in the workflow of global optimization since it can produce more low-energy configurations close to the DFT global minimum and reduce the number of DFT calculations, however, the improved accuracy generally comes with an increased computational cost. Though minimization is fast for a single structure and can be made parallel, the computational cost can become intensive for complex systems even by MLIPs, where numerous structures are calculated.
The benchmark on the MD of supported nanoparticles shows that the uMLIPs can produce consistent results of MSDCu and RDF in 20-ps-long simulations compared to AIMD. When it comes to long-time MD simulations, it becomes challenging in two aspects: the lack of DFT reference and the computational cost. The property predicted by MLIPs in the long-time limit can only be cross-validated by different models in the absence of AIMD. MACE-OMAT and MatterSim-v1.0.0-1M exhibit qualitative trends in MSDCu up to 100 ps, which are similar to our DP-UniAlCu model and the MACE-UniAlCu trained on our dataset. To obtain more quantitative results, longer MD simulations are required but not feasible for uMLIPs. Our DP-UniAlCu model is about two orders of magnitude faster than MACE-OMAT and Mattersim-v1.0.0-1M, which makes it a good candidate for long-time MD simulations. Currently, the best tradeoff in accuracy and efficiency is still obtained within the DP architecture. The accuracy of DP can be improved by message passing as demonstrated with DP-MPGao2024PCCP and the efficiency is not degraded as much as switching to uMLIPs. Nevertheless, the uMLIPs can be an alternative method to generate configurations through MD simulations for the initial dataset in the MLIP development instead of the perturbation of a few given structures.Zhang2019PRM For example, distillation is a promising strategy to alleviate the computational performance issue of uMLIPs by training a cheaper model on the configurations sampled and labelled by a uMLIP.
The emergence of uMLIPs provides a good foundation to develop domain-specific MLIPs for complex systems, however, the challenges still exist in reducing the computational cost and integrating fine-tuning and distillation in the current framework of active learning. We expect the benchmarks in this work can provide valuable insights into the applications of uMLIPs for a better understanding of supported nanoparticle catalysts.
@@ 4 Methods
The DFT calculations used in binding energy computation and ab initio molecular dynamics are performed with the Vienna Ab initio Simulation Package (VASP).Kresse1996ComputMaterSci, Kresse1996PRB The Perdew-Burke-Ernzerhof (PBE) approximation for the exchange-correlation functionalPerdew1996PRL is used in our UniAlCu dataset and in calculating the binding energies of supported nanoparticles. The MACE-MP, MACE-OMAT, MatterSim-v1.0.0-1M, and MatterSim-v1.0.0-5M are trained on the PBE-based datasets with slightly different parameters. The detailed comparison of input settings can be found in Section S2. The MACE-UniAlCu model is trained on our UniAlCu dataset using the same setting as MACE-OMAT, which is discussed in Section S3.
The simulations by uMLIPs are performed with either LAMMPSThompson2022ComputPhysCommun or ASELarsen2017JPhysCondensMatter through interfaces implemented in the GDPy package.Xu2024GDPy The global optimization is based on the genetic algorithm implemented in the GDPy package,Lee2022JPCC, Xu2024GDPy where the crossover used a cut-and-splice algorithmDeaven1995PRL, Vilhelmsen2012PRL and the mutations include the rattle and mirror operations with equal probabilities. The population size is set to 50 and the number of generations is set to 20. The configurations in the first generation are randomly generated, and those in the following generations are produced by the crossover and mutation methods.
@@
where $E_{\mathrm{Cu}_{N}/surf}$ is the potential energy of a Cu nanoparticle on an Al2O3 surface, $E_{surf}$ is the potential energy of a pristine Al2O3 surface, and $E_{\mathrm{Cu}_{N}}$ is the potential energy of the global minimum configuration of CuN in the gas phase.
Two ranking metrics are used to quantify the quality of an MLIP-generated low-energy ensemble. The rank based overlap (RBO)Webber2010ACMTOIS is defined as
@@
[…]
@@
The mean-squared displacements of Cu atoms (MSDCu) is given by
$MSD(t;r)={<\frac{1}{N}\sum_{i=1}^{N}\lVert r(t+\tau)-r(\tau)\rVert^{2}>}_{\tau},$ (4)
@@
{acknowledgement}
@@
The funding for this project was provided by Shell International Exploration and Production Inc., USA. The calculations were performed largely using the Princeton Research Computing resources at Princeton University. The authors also acknowledge the computing resources from the National Energy Research Scientific Computing Center (NERSC) operated under Contract No. DE-AC0205CH11231 using NERSC award ERCAP0021510. The software used in the present work has been developed in the Computational Chemical Science Center “Chemistry in Solution and at Interfaces (CSI)” funded by the USA Department of Energy under award DE-SC0019394. RC was partially supported by this grant, and JX benefited of scientific discussions with the members of CSI.
{suppinfo}
7 figures, 2 tables, notes and references
Figure 1: Several MLIPs are examined for the binding energies of Cu1-21 on three Al2O3 surfaces: (a) $\gamma$-Al2O3(100) (b) $\gamma$-Al2O3(110) (c) $\alpha$-Al2O3(0001) (d) violin plots of errors in binding energy. DP-UniAlCu and MACE-UniAlCu have smaller spreads on the binding energy errors since they are specifically trained on these systems. Among uMLIPs, MACE-OMAT and MatterSim-v1.0.0-5M show similar accuracy as DP-UniAlCu, which reproduce the relative binding strength across three surfaces. Degraded performance is observed for MACE-MP and MatterSim-v1.0.0-1M as the former one is trained on a smaller dataset and the latter one is a model with smaller parameters. DPA2-MPTrj shows large energy differences compared to DFT and DPA3-OMAT has reduced errors due to its improved architecture.
Table 1: Energetics of Global Minimum Configurations (units in eV)
Model DP-UniAlCu MACE-OMAT MatterSim-v1.0.0-1M
Surface Size $avg(\Delta E_{b})\downarrow$ $std(\Delta E_{b})\downarrow$ $avg(\Delta E_{b})\downarrow$ $std(\Delta E_{b})\downarrow$ $avg(\Delta E_{b})\downarrow$ $std(\Delta E_{b})\downarrow$
$\gamma$-Al2O3(100) Cu27 -0.56 0.27 -1.95 0.28 -9.67 0.48
Cu38 -0.99 0.22 -2.59 0.54 -13.74 0.55
Cu47 -1.21 0.32 -2.69 0.25 -16.83 0.40
Cu55 -1.33 0.26 -0.80 0.19 1.48 0.35
$\gamma$-Al2O3(110) Cu27 -0.42 0.30 -1.53 0.38 -8.00 0.42
Cu38 -0.70 0.39 -1.85 0.45 -11.91 0.72
Cu47 -0.96 0.55 -2.00 0.55 -14.74 0.93
Cu55 -1.11 0.45 -0.15 0.16 3.34 0.55
$\alpha$-Al2O3(0001) Cu27 -0.27 0.20 -0.94 0.20 1.72 0.49
Cu38 -0.10 0.23 -0.94 0.22 2.36 0.74
Cu47 0.07 0.47 -1.05 0.23 2.17 0.31
Cu55 -0.51 0.47 -1.61 0.19 2.25 1.08
Table 2: Ranking Metrics on Global Minimum Configurations
Surface Size
[…]
Kendall $\tau\uparrow$
DP-UniAlCu MACE-OMAT MatterSim-v1.0.0-1M
$\gamma$-Al2O3(100) Cu27 q 0.36 0.29 0.02
$\tau$ 0.58 0.26 0.05
Cu38 q 0.41 0.01 0.02
$\tau$ 0.63 -0.42 0.07
Cu47 q 0.70 0.21 0.79
$\tau$ 0.51 0.70 0.46
Cu55 q 0.44 0.30 0.26
$\tau$ 0.48 0.48 0.39
$\gamma$-Al2O3(110) Cu27 q 0.37 0.13 0.06
$\tau$ 0.60 0.51 0.46
Cu38 q 0.75 0.28 0.52
$\tau$ 0.59 0.59 0.56
Cu47 q 0.15 0.02 0.01
$\tau$ 0.21 -0.27 -0.14
Cu55 q 0.46 0.33 0.49
$\tau$ 0.26 0.56 -0.01
$\alpha$-Al2O3(0001) Cu27 q 0.14 0.28 0.01
$\tau$ 0.50 0.61 -0.20
Cu48 q 0.54 0.12 0.05
$\tau$ 0.42 0.12 0.05
Cu47 q 0.43 0.51 0.46
$\tau$ 0.27 0.81 0.61
Cu55 q 0.43 0.25 0.01
$\tau$ -0.06 0.43 -0.21
Figure 2: The low-energy configurations of Cu27 and Cu38 on $\gamma$-Al2O3(100), $\gamma$-Al2O3(110), and $\alpha$-Al2O3(0001) by DFT, DP-UniAlCu (blue), MACE-OMAT (green), and MatterSim-v1.0.0-1M (orange). The DFT configurations are made up of structures obtained from MLIP-based global optimization. Each MLIP provides 50 structures for a given combination of the nanoparticle size and the supporting surface, forming a low-energy ensemble of 150 structures for each system. The arrows indicate the positions of the DFT-minimized global minimum configurations in spectra.
Figure 3: The low-energy configurations of Cu47 and Cu55 on $\gamma$-Al2O3(100), $\gamma$-Al2O3(110), and $\alpha$-Al2O3(0001) by DFT, DP-UniAlCu (blue), MACE-OMAT (green), and MatterSim-v1.0.0-1M (orange). The DFT configurations are made up of structures obtained from MLIP-based global optimization. Each MLIP provides 50 structures for a given combination of the nanoparticle size and the supporting surface, forming a low-energy ensemble of 150 structures for each system. The arrows indicate the positions of the DFT-minimized global minimum configurations in spectra.
Figure 4: For each MLIP, 20 MD simulations starting from the same global minimum configuration of Cu13 but initialized from different velocities by Maxwell-Boltzmann distribution at 800 K are performed for 20 ps with a timestep of 2 fs. For AIMD, only one trajectory is performed with the same settings. (a)-(c) The mean square displacement of Cu atoms. The vertical bars are the standard deviations of 20 MD simulations by MLIPs. (d)-(f) The radial distribution functions of Cu-Al. (g)-(i) The radial distribution functions of Cu-O.
Figure 5: (a)-(c) The mean square displacement of Cu atoms for Cu13 on three Al2O3 surfaces at 800 K. (d) The computational performance of DP-UniAlCu, DPA3-OMAT, MACE-OMAT, MACE-UniAlCu, and MatterSim-v1.0.0-1M models on the simulation of an FCC Cu bulk with 2048 atoms. The “katoms*step/s” is defined as the number of thousands of atoms processed per MD step per second. The inset panels show the initial structures of MD simulations.

### v2

@@ 2 Results
We first tested the mean-squared displacements of Cu atoms (MSDCu, defined in Equation 4) of Cu13 on three Al2O3 surfaces with 20-ps MD simulations, where ab initio MD (AIMD) trajectories are available for reference. The DP-UniAlCu, MACE-UniAlCu, MACE-OMAT, MatterSim-v1.0.0-1M, MatterSim-v1.0.0-5M, and DPA3-OMAT are tested. For each MLIP, 20 MD simulations starting from the same minimum energy configuration but initialized from different velocities by Maxwell-Boltzmann distribution at 800 K are performed for 20 ps with a timestep of 2 fs. For AIMD, only one trajectory is performed with the same settings. The window-averaged MSDCu up to 5 ps are shown in Figure 4a-c. The vertical bars are the standard deviations of 20 MD simulations by MLIPs. The model accuracy can be examined by checking whether the AIMD results fall within the standard deviations of MLIPs. For $\gamma$-Al2O3(100), all models except for MACE-OMAT and DPA3-OMAT give MSDCu results that are close to the AIMD results, which exhibit smaller MSDCu values. For $\gamma$-Al2O3(110), all models give consistent results with AIMD. For $\alpha$-Al2O3(0001), all models except for two MatterSim models produce MSDCu results closer to AIMD, which show larger MSDCu values. The radial distribution functions (RDFs) of Cu-Al and Cu-O pairs are shown in Figure 4d-f and Figure 4g-i, respectively, which are averaged over snapshots every 0.1 ps from 0.2 ps to 20 ps. The RDFs of Cu-Al and Cu-O by DP-UniAlCu, MACE-UniAlCu and MACE-OMAT are in good agreement with the AIMD results, while other models give larger deviations. The RDF of Cu-Cu can be well reproduced by all models (Figure S8). In addition to the above models, we also tested DPA2-MPTrj and DPA2-OC2M. However, MD simulations with these models were not stable and collapsed within a few ps at 800 K, consistent with the large force errors on the UniAlCu dataset shown in Table S6.
Since the 20-ps-long MD simulations are too short to observe any diffusion of the entire nanoparticle, we performed 1-ns-long MD simulations for Cu13 on three $\alpha$-Al2O3(0001) surfaces with DP-UniAlCu, MACE-UniAlCu, MACE-OMAT, MatterSim-v1.0.0-1M, and DPA3-OMAT to investigate the long-time dynamic behavior. Five independent MD simulations are performed for each MLIP starting from the same minimum energy configuration but initialized from different velocities by Maxwell-Boltzmann distribution at 800 K. The results of MSDCu up to 100 ps for Cu13 on three Al2O3 surfaces are shown in Figure 5a-c. The vertical bars are the standard deviations of five MD simulations by MLIPs. Since ns-long MD simulations are not readily accessible to AIMD, the benchmark here only served as a cross-validation across different MLIPs by comparing results against our DP-UniAlCu model. All the above MLIPs predict qualitatively consistent mobility trends for the Cu atoms in the nanoparticles supported on the three Al2O3 surfaces. The mobility of Cu13 on these three surfaces was discussed in previous work,29 where it was found that MSDCu on $\gamma$-Al2O3(100) and $\alpha$-Al2O3(0001) are comparable and larger than on $\gamma$-Al2O3(110). MatterSim-v1.0.0-1M gives slightly larger MSDCu for Cu13 on $\gamma$-Al2O3(100) and $\gamma$-Al2O3(110), and much larger MSDCu for $\alpha$-Al2O3(0001) compared to DP-UniAlCu. MACE-OMAT gives almost identical MSDCu for Cu13 on $\gamma$-Al2O3(110) as DP-UniAlCu while more deviations between them are observed for Cu13 on $\gamma$-Al2O3(100) and $\alpha$-Al2O3(0001), where MACE-OMAT gives smaller MSDCu values. By using the MACE-UniAlCu model, which is trained on the same dataset as DP-UniAlCu, the MSDCu results for $\gamma$-Al2O3(100) and $\alpha$-Al2O2(0001) become closer to those by DP-UniAlCu, however, a smaller MSDCu is observed for $\gamma$-Al2O3(110). DPA3-OMAT exhibits a consistent mobility trend as DP-UniAlCu but gives smaller MSDCu values for all three Al2O3 surfaces. To perform a more quantitative comparison, more simulations with a longer time are necessary, which can be challenging for uMLIPs. The computational efficiency is shown in Figure 5d, which is taken from the 100-ps simulation of an FCC Cu bulk with 2048 atoms. All models are performed on a single NVIDIA 80G A100 GPU with LAMMPS23 except for MatterSim-v1.0.0-1M with ASE15. DP-UniAlCu is roughly two orders of magnitude faster than the DPA3-OMAT and MACE-based models and more than one order of magnitude faster than MatterSim-v1.0.0-1M. The uMLIPs and domain-specific MLIPs serve different purposes. Domain-specific applications that require large size simulations in space and/or time may be only accessible with domain-specific models with available computational resources.
3 Discussion
@@
In this work, we tested several uMLIPs on supported Cu nanoparticles on Al2O3 surfaces and found that they can be very useful in simulating supported nanoparticles even without any fine-tuning, including the global optimization of low-energy structures and the MD at finite temperature.
@@
The uMLIPs examined in this work can produce reasonable results on the binding energies of small supported nanoparticles (Cu1-21) reported in our previous study,29 where MACE-OMAT shows an accuracy comparable to our domain-specific DP-UniAlCu model. In the global optimization of large supported nanoparticles (Cu27, Cu38, Cu47, and Cu55), our domain-specific DP-UniAlCu model more frequently locates low-energy structures than MACE-OMAT and Mattersim-v1.0.0-1M. Interestingly, Mattersim-v1.0.0-1M, which shows larger deviations in binding energies, can find even more stable configurations than the other two models in some systems, indicating its exploration ability in the configuration space is not solely characterized by the energetic metrics on a preset dataset. Alongside the exploration ability of MatterSim-v1.0.0-1M, the non-uniform accuracy across different systems is observed, for instance, the binding energy accuracy in one nanoparticle size is better than another size on the same Al2O3 surface. This non-uniform accuracy may be due to the imbalance in the number of configurations from different systems, where the initial dataset is made up of configurations from one specific system and the configurations of other systems are accumulated in active learning. In terms of global optimization, since the system size of interest is usually accessible to DFT, the MLIP-generated configurations can still be verified by DFT to find the low-energy structures. The model’s stability in generating diverse configurations in a large region of the configuration space is more important than the accuracy of the energy, as it provides more candidates for DFT verification. A model with better accuracy can improve the performance in the workflow of global optimization since it can produce more low-energy structures close to DFT and reduce the number of DFT calculations, however, the improved accuracy generally comes with an increased computational cost. Though minimization is fast for a single structure and can be made parallel, the computational cost can become intensive for complex systems even by MLIPs, where numerous structures are calculated.
The MD benchmark for supported nanoparticles shows that uMLIPs can produce results consistent with AIMD for MSDCu and RDF in simulations having a 20 ps timescale. When larger scale MD simulations are required, benchmarking becomes challenging in two aspects: the lack of DFT reference and the computational cost. The properties predicted by the MLIPs in the long-time limit can only be cross-validated within the models in the absence of AIMD. MACE-OMAT and MatterSim-v1.0.0-1M exhibit qualitative trends for MSDCu up to 100 ps similar to the DP-UniAlCu and MACE-UniAlCu models, but this is still too short to reliably estimate MSDCu. Longer MD simulations are necessary for that, but they may not be feasible with uMLIPs. DP-UniAlCu is about two orders of magnitude faster than MACE-OMAT and Mattersim-v1.0.0-1M. The latter models are not only universal, but also have better accuracy for energy and forces due to their more advanced architectures. However, the energy and force accuracy of the standard DP model can be improved with marginal additional cost by adopting message passing and a slightly higher level of equivariance  9. uMLIPs can be very useful in the development of system specific MLIPs particularly in the initial stages of construction of the training dataset, when this is typically not sufficiently structurally diverse.35 For this purpose, DFT labeling can be replaced by uMLIP labeling, gaining both structural diversity and computational cost. DFT labeling can then be utilized more rarely only in the final stages of the construction of the data set with a substantial gain in computational cost. Ideally, a uMLIP could be distilled into an efficient model specific MLIP without any DFT labeling. Our work shows that this procedure can be problematic because the accuracy of the uMLIPs may be not uniform, and they can occasionally generate unphysical structures. In this context, a possible way to be explored would require using uncertainty quantification techniques, such as discussed in Refs. 3, 17, as a diagnostic step to assess whether a given uMLIP reliably describes the relevant configuration space before distillation. The benchmarks discussed in this work should provide useful insight on how uMLIPs could be used in application studies of supported nanoparticle catalysts.
@@ 4 Methods
4.1 Density Functional Theory and Machine Learning Interatomic Potentials
The DFT calculations used in binding energy computation and ab initio molecular dynamics are performed with the Vienna Ab initio Simulation Package (VASP).13, 14 The Perdew-Burke-Ernzerhof (PBE) approximation for the exchange-correlation functional21 is used in our UniAlCu dataset and in calculating the binding energies of supported nanoparticles. The MACE-MP, MACE-OMAT, MatterSim-v1.0.0-1M, and MatterSim-v1.0.0-5M are trained on the PBE-based datasets with slightly different parameters. The detailed comparison of input settings can be found in Section S2. The MACE-UniAlCu model is trained on our UniAlCu dataset using the same setting as MACE-OMAT, which is discussed in Section S3.
4.2 Molecular Dynamics and Global Optimization
The simulations by uMLIPs are performed with either LAMMPS23 or ASE15 through interfaces implemented in the GDPy package.31 The global optimization is based on the genetic algorithm implemented in the GDPy package,16, 31 where the crossover used a cut-and-splice algorithm8, 25 and the mutations include the rattle and mirror operations with equal probabilities. The population size is set to 50 and the number of generations is set to 20. The configurations in the first generation are randomly generated, and those in the following generations are produced by the crossover and mutation methods.
4.3 Benchmarking Metrics
@@
where $E_{\mathrm{Cu}_{N}/surf}$ is the potential energy of a Cu nanoparticle on an Al2O3 surface, $E_{surf}$ is the potential energy of a pristine Al2O3 surface, and $E_{\mathrm{Cu}_{N}}$ is the potential energy of the minimum energy configuration of CuN in the gas phase.
Two ranking metrics are used to quantify the quality of MLIP-generated low-energy structures. The RBO26 is defined as
@@
where $S$ and $T$ are two lists of binding energies ranked from low to high, $d$ is the length of a sub-list, and $p$ is a parameter controlling the top-weightness and chosen to be 0.94 when 95% of the weight is contributed by the top fifty energies. The Kendall’s $\tau$12 is defined as
@@
The MSDCu is given by
$\mathrm{MSD}(t;r)={<\frac{1}{N}\sum_{i=1}^{N}\lVert r(t+\tau)-r(\tau)\rVert^{2}>}_{\tau},$ (4)
@@
5 Data Availability
The UniAlCu dataset is available online,30 with information given in the respective publication.29
6 Acknowledgements
@@
The funding for this project was provided by Shell International Exploration and Production Inc., USA. The calculations were performed largely using the Princeton Research Computing resources at Princeton University. The authors also acknowledge the computing resources from the National Energy Research Scientific Computing Center (NERSC) operated under Contract No. DE-AC0205CH11231 using NERSC award ERCAP0021510. The software used in the present work has been developed in the Computational Chemical Science Center “Chemistry in Solution and at Interfaces (CSI)” funded by the USA Department of Energy under award DE-SC0019394. R.C. was partially supported by this grant.
7 Supplementary Information
8 figures, 6 tables, notes and references
8 Author contributions
J.X., A.P., A.D.P., S.S., and R.C. conceived the idea of this work and designed the benchmark. J.X. performed the calculations and wrote the original draft. J.X., A.P., S.S., and R.C. revised the manuscript. J.X., A.P., A.D.P., S.S., D.H., and R.C. contributed to the discussions and reviewed the manuscript.
9 Competing interests
The authors declare no competing financial or non-financial interests.